find the slope for 9x-2y=7

Answer by NAMath(12)

Arrange the given line according to the slop-intercept form
9x-2y=7
-2y=-9x+7
Divide by -2 both sides
-2y/-2=(-9x+7)/-2
y=((-9x)/-2)+(7/-2)
y=(9/2)x-7/2
Compare above equation with the equation of line slope-intercept form
y=mx+b
y=(9/2)x-7/2
m=9/2 and b=-7/2
Slope of the given line m = 9/2